# Two algebraic proofs of the transcendence of $\mathrm{e}$ based on formal power series

### v1

@@ 3 The proof by formal integration
We present the second semiformal proof of the transcendence of $\mathrm{e}$. It is due to this author and first appeared in early versions of [4]. The idea is simply to translate Hilbert’s proof, which we described at the beginning, into the language of formal power series. At the end of the section we explain why the proof is semiformal and why conditions (A) and (B) are satisfied.
We work in the domains of formal power series $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$. We introduce additional operations on $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$ besides those we reviewed in Section 1. For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
@@
This is a formal operation. Clearly, for every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
@@
If $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ then $f(0x)=a_{0}$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$.
In Section 1 we defined the derivative $f^{\prime}$ and the primitive $\int f$ of a formal power series $f\in\mathbb{R}[[x]]$. In the next proposition, whose straightforward proof we omit, we collect properties of these operations.
@@
The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent, and $f(x+b)\in\mathbb{R}\{x\}$. It is easy to see that $f(x+0)=f(x)$ and that if $f(x)\in\mathbb{R}[x]$ then $f(x+b)$ is the usual substitution in a polynomial. In the next two propositions we show that semiformal shifts behave in the expected way.
@@
Let $f(x)\in\mathbb{R}\{x\}$, $c,d\in\mathbb{R}$ and $g(x)=f(x+c)$. Then
g(x+d)=f(x+(c+d)).g(x+d)=f(x+(c+d))\;.
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $c$, $d$ and $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ be as stated. Then
$b_{n}=\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,g(x+d)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}b_{m}d^{m-n}\big)x^{n}\,.$}
Thus the coefficient of $x^{n}$ in $g(x+d)$ equals
${\textstyle\sum_{m\geq n}\binom{m}{n}b_{m}d^{m-n}=\sum_{m\geq n}\binom{m}{n}\big(\sum_{l\geq m}\binom{l}{m}a_{l}c^{l-m}\big)d^{m-n}=}$
@@
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{j=0}^{l-n}\binom{l-n}{j}c^{j}d^{l-n-j}=\sum_{l\geq n}\binom{l}{n}a_{l}(c+d)^{l-n}\,,}$
which is the coefficient of $x^{n}$ in $f(x+(c+d))$. In the third line, we changed the order of summation. This is allowed because $f(x)$ is convergent. In the fourth line, we used the binomial theorem. $\Box$
@@
Let $f,g\in\mathbb{R}\{x\}$ and $c\in\mathbb{R}$. Then
@@
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ and $c$ be as stated. The coefficient of $x^{n}$ in the formal power series $f(x+c)$, respectively $g(x+c)$, equals $\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}$, respectively $\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}$. Thus the coefficient of $x^{n}$ in the product $f(x+c)\cdot g(x+c)$ equals
${\textstyle\sum_{k=0}^{n}\big(\sum_{m\geq k}\binom{m}{k}a_{m}c^{m-k}\cdot\sum_{l\geq n-k}\binom{l}{n-k}b_{l}c^{l+k-n}\big)=}$
@@
${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\binom{m+l}{n}a_{m}b_{l}c^{l+m-n}=\sum_{j\geq n}\binom{j}{n}\big(\sum_{m=0}^{j}a_{m}b_{j-m}\big)c^{j-n}\,,}$
which is the coefficient of $x^{n}$ in $(f\cdot g)(x+c)$. In the third line, we changed the order of summation. This is allowed since $f(x)$ is convergent. In the fourth line, we used Vandermonde’s identity ([6]). $\Box$
We prove a semiformal exponential identity.
Theorem 3.6
Let $a,b\in\mathbb{R}$ with $a\neq 0$ and $f(x)=\mathrm{Exp}\big(x+\frac{b}{a}\big)$. Then
$f(ax)=\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$
Proof. Let $a$, $b$ and $f(x)$ be as stated. Then, indeed,
$f(ax)$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}(a^{m}/m!)(b/a)^{m-n}\big)x^{n}}$
$=$ ${\textstyle\sum_{n\geq 0}\sum_{m\geq n}\frac{b^{m-n}}{(m-n)!}(a^{n}/n!)x^{n}=\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)\frac{a^{n}}{n!}x^{n}}$
@@
We introduce the Newton integral of convergent formal power series.
@@
The integral is always defined by item 3 of Proposition 3.1 because $\int f\in\mathbb{R}\{x\}$. The Newton integral is a semiformal operation because it uses two values of a formal power series. We establish for it three integral transformations: linear combination, additive combination, and shift of the integration interval. The first two are quite straightforward.
@@
In the third line, we used the identity $\binom{m+1}{n+1}=\frac{m+1}{n+1}\binom{m}{n}$ and we exchanged the order of summation. This is allowed since the involved series absolutely converge. In the fourth line, we add and subtract the term $b^{m+1}$ in the two inner sums. In the fifth line, we use the binomial theorem. $\Box$
We introduce the improper Newton integral of formal power series.
@@
Let $u\in\mathbb{R}$ and $f(x)\in\mathbb{R}\{x\}$. If for every real sequence $(b_{n})$ going to $\infty$ the limit
@@
We call this integral the improper Newton integral from $u$ to $\infty$ of the formal power series $f(x)$.
@@
We obtain for the improper integral the same three integral transformations as before.
@@
The third transformation is again less immediate.
@@
Let $f\in\mathbb{R}\{x\}$ and $u,b\in\mathbb{R}$. The equality
@@
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $u$ and $b$ be as stated. Let $(c_{n})$ be any real sequence converging to $\infty$, and let $L$ and $R$ be the integrals on the left-hand side and the right-hand side, respectively. It suffices to show the implication $\exists\,R\Rightarrow R=L$; the opposite implication then follows. Indeed, we suppose that integral $L$ exists and write it as
$L=\int_{u+b-b}^{+\infty}f(x+b)\,.$}
We set $g(x)=f(x+b)$, use the assumed implication with b-b and get that
@@
Thus, the implication $\exists\,L\Rightarrow L=R$ holds as well.
We suppose that integral $R$ exists. We replace the test sequence $(c_{n})$ with $(c_{n}+b)$ and get
$R$ $=$ $\lim_{m\to\infty}{\textstyle\big(\sum_{n\geq 0}a_{n}\frac{(c_{m}+b)^{n+1}}{n+1}\big)-\sum_{n\geq 0}a_{n}\frac{(u+b)^{n+1}}{n+1}}$
$=$ $\lim_{m\to\infty}{\textstyle\big(\sum_{l\geq 0}\sum_{n\geq l}\binom{n+1}{l+1}a_{n}b^{n-l}\frac{c_{m}^{l+1}}{n+1}\big)-\sum_{l\geq 0}\sum_{n\geq l}\binom{n+1}{l+1}a_{n}b^{n-l}\frac{u^{l+1}}{n+1}}$
$=$ $\lim_{m\to\infty}{\textstyle\big(\sum_{l\geq 0}\sum_{n\geq l}\binom{n}{l}a_{n}b^{n-l}\frac{c_{m}^{l+1}}{l+1}\big)-\sum_{l\geq 0}\sum_{n\geq l}\binom{n}{l}a_{n}b^{n-l}\frac{u^{l+1}}{l+1}}$
$=$ $\lim_{m\to\infty}{\textstyle\big(\int f(x+b)\big)(c_{m})}-{\textstyle\big(\int f(x+b)\big)(u)}=L\,.$
In the second line, we used the binomial theorem, and we changed the order of summation, which is allowed by the convergence of $f(x)$. Also, we canceled the summands corresponding to $l+1=0$. In the third line, we used the transformation $\binom{l+1}{n+1}=\frac{l+1}{n+1}\binom{l}{n}$. In the fourth line, we use the definition of formal primitive and Definition 3.3. $\Box$
We need bounds on integrals
$\int_{0}^{i}p(x)\mathrm{Exp}(-x)$}
where $i\in\mathbb{N}_{0}$, $p(x)$ is an integral polynomial and $\mathrm{Exp}(-x)=\mathrm{Exp}((-1)x)$. In the semiformal approach, this appears problematic because one deduces the standard LM bound by considering all values of the integrand on $[0,i]$ or $0,i]\cap\mathbb{Q}$. Instead, we use the next substitute LM bound, which can be proven by means of just two values of the primitive.
@@
$\text{$\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!}x^{n+k}$ has the primitive $\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!(n+k+1)}x^{n+k+1}$}\,.$}
Thus the Newton integral is in absolute value at most $\sum_{n=0}^{\infty}\frac{1}{n!}i^{n+k+1}+0=i^{k+1}\mathrm{e}^{i}$. $\Box$
It is important that the bound is exponential and not factorial in $k$.
The key step in Hilbert’s proof of the transcendence of $\mathrm{e}$ was Euler’s identity. Now we prove it in the setting of the improper Newton integral of formal power series.
Theorem 3.16
@@
$I_{k}=\int_{0}^{+\infty}x^{k}\cdot\mathrm{Exp}(-x)=\int_{0}^{+\infty}\sum_{n=k}^{\infty}\frac{(-1)^{n-k}}{(n-k)!}x^{n}=k!\,.$}
Proof. We proceed by induction on $k$. Let $(a_{n})$ be any real sequence going to $\infty$. For $k=0$ we use that
$\int\mathrm{Exp}(-x)=\sum_{n\geq 0}\frac{(-1)^{n}}{(n+1)!}x^{n+1}=1-\mathrm{Exp}(-x)\,.$}
Hence
@@
For $k\in\mathbb{N}_{0}$, let $F_{k}=x^{k}\cdot\mathrm{Exp}(-x)$. Let $k\in\mathbb{N}$. By item 1 of Proposition 3.1 and the derivative of $\mathrm{Exp}(-x)$,
$F_{k}^{\prime}=\big(x^{k}\cdot\mathrm{Exp}(-x)\big)^{\prime}=kx^{k-1}\cdot\mathrm{Exp}(-x)-x^{k}\cdot\mathrm{Exp}(-x)=k\cdot F_{k-1}-F_{k}\,.$
By item 4 of Proposition 3.1, $\int F_{k}^{\prime}=F_{k}$. Thus by Proposition 3.12 and induction,
@@
$=$ $k\cdot I_{k-1}-\lim_{n\to\infty}F_{k}(a_{n})+F_{k}(0)$
$=$ $k\cdot(k-1)!-0+0=k!\;.$
@@
where $a_{i}\in\mathbb{Z}$, $n\in\mathbb{N}$, and $a_{0}\neq 0$. We multiply this equality by the improper Newton integral (of formal power series)
@@
where $p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}$ (as in Section 1). This integral exists by Proposition 3.12 and Theorem 3.16. Using Proposition 3.13 we get that
@@
for every $r\in\mathbb{N}$. We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded.
The determination of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that we calculate semiformally. We define abbreviations
$\mathrm{Exp}(-(x-i))$ $=$ $f((-1)x),\ f(x)=\mathrm{Exp}(x+(-i))\,\text{ and}$
$\mathrm{Exp}(-((x-i)+i))$ $=$ $f((-1)x),\ f(x)=g(x+i),\ g(x)=\mathrm{Exp}(x+(-i))\,.$
Then
@@
Recall that $x$ is a formal variable. Using Proposition 3.12, Theorem 3.16, and the definition of the polynomial $p_{r}(x)$, we obtain the properties of $B_{r}$ as before.
It remains to bound the numbers
@@
has degree less than $(n+1)(r+1)$ and each coefficient of it is in absolute value at most $l^{r+1}$ where $l\in\mathbb{N}$ depends only on $n$. By Propositions 3.8 and 3.15 each of the above $n+1$ Newton integrals in $A_{r}$ is in absolute value at most
@@
Thus $|A_{r}|\leq c^{r}$ for every $r\in\mathbb{N}$ and some constant $c\geq 1$. We get the same contradiction as in Section 1. $\Box$
The previous proof is quasiformal and not formal because of Definitions 3.3 and 3.11. It clearly satisfies conditions (A) and (B).

### v3

@@ 3 The proof by formal integration
We present the second AB &\& SF proof of the transcendence of $\mathrm{e}$. It is due to this author and first appeared in early versions of [7]. The idea is to translate Hilbert’s proof, which we described at the beginning, into the language of formal power series.
We work in the domains of formal power series $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$. We introduce further operations on $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$ besides those we mentioned in Section 1. For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
@@
The operation $f(x)\mapsto f(ax)$, which we call the constant inner product, is formal. Clearly, for every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
@@
If $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ then $f(0x)=a_{0}$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$. By f(x)f(-x) we mean f((1)x)f((-1)x).
In the next proposition, whose straightforward proof we omit, we collect properties of the operations $f\mapsto f^{\prime}$ and $f\mapsto\int f$.
@@
The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent. It is easy to show that $f(x+b)\in\mathbb{R}\{x\}$. Clearly, $f(x+0)=f(x)$ and if $f(x)\in\mathbb{R}[x]$ then $f(x)\mapsto f(x+b)$ is the usual polynomial substitution. We combine the operations of constant inner product and semiformal shift by means of brackets. For example, $f((x+a)+b)$ is $g(x+b)$ where $g(x)=f(x+a)$ or $f((ax)+b)$ is $g(x+b)$ where $g(x)=f(ax)$ or $f(a(x+b))$ is $g(ax)$ where $g(x)=f(x+b)$ and so on.
In the next two propositions, we show that semiformal shifts behave in desirable ways.
@@
Let $f(x)\in\mathbb{R}\{x\}$ and $c,d\in\mathbb{R}$. Then
f((x+c)+d)=f(x+(c+d)).f((x+c)+d)=f(x+(c+d))\,.
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ and $c,d$ be as stated, and let $f(x+c)=\sum_{n\geq 0}b_{n}x^{n}$. Then
$b_{n}=\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,f((x+c)+d)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}b_{m}d^{m-n}\big)x^{n}\,.$}
Thus
${\textstyle[x^{n}]f((x+c)+d)=\sum_{m\geq n}\binom{m}{n}\big(\sum_{l\geq m}\binom{l}{m}a_{l}c^{l-m}\big)d^{m-n}=}$
@@
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{j=0}^{l-n}\binom{l-n}{j}c^{j}d^{l-n-j}=\sum_{l\geq n}\binom{l}{n}a_{l}(c+d)^{l-n}=}$
$=[x^{n}]f(x+(c+d))\,.$
In the third line, we changed the order of summation. This is allowed because $f(x)$ is convergent. In the fourth line, we used the binomial theorem. $\Box$
@@
Let $f(x),g(x)\in\mathbb{R}\{x\}$ and $c\in\mathbb{R}$. Then
@@
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ and $c$ be as stated. The coefficient of $x^{n}$ in the formal power series $f(x+c)$, respectively $g(x+c)$, equals $\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}$, respectively $\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}$. Thus
${\textstyle[x^{n}]f(x+c)g(x+c)=\sum_{k=0}^{n}\big(\sum_{m\geq k}\binom{m}{k}a_{m}c^{m-k}\sum_{l\geq n-k}\binom{l}{n-k}b_{l}c^{l+k-n}\big)}$
@@
${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\binom{m+l}{n}a_{m}b_{l}c^{l+m-n}=\sum_{j\geq n}\binom{j}{n}\big(\sum_{m=0}^{j}a_{m}b_{j-m}\big)c^{j-n}=}$
$=[x^{n}](fg)(x+c)\,.$
In the third line, we changed the order of summation. This is allowed since $f(x)$ is convergent. In the fourth line, we used Vandermonde’s identity ([9]). $\Box$
The following semiformal exponential identity is somewhat counter-intuitive (at first we got it wrong).
Proposition 3.6
Let $a,b\in\mathbb{R}$ with $a\neq 0$. Then
$\mathrm{Exp}\big((ax)+\frac{b}{a}\big)=\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$}
Proof. Let $a$, $b$ be as stated, and let $f(x)=\mathrm{Exp}(ax)$. Then
${\textstyle f\big(x+\frac{b}{a}\big)}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}(a^{m}/m!)(b/a)^{m-n}\big)x^{n}}$
$=$ ${\textstyle\sum_{n\geq 0}\sum_{m\geq n}\frac{b^{m-n}}{(m-n)!}(a^{n}/n!)x^{n}=\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)(a^{n}/n!)x^{n}}$
@@
We will use this identity only with a=1a=-1.
We introduce Newton integrals of convergent formal power series.
@@
By item 3 of Proposition 3.1, $\int f\in\mathbb{R}\{x\}$ and the integral is always defined. The Newton integral is a semiformal operation because it uses two values of a formal power series. We prove three integral transformations for it: linear combination, additive decomposition, and shift of the integration interval. The first two transformations are straightforward.
@@
In the third line, we used the identity $\binom{m+1}{n+1}=\frac{m+1}{n+1}\binom{m}{n}$ and we exchanged the order of summation. This is allowed since the involved series absolutely converge. In the fourth line, we add
$0=\sum_{m\geq 0}\frac{a_{m}}{m+1}\binom{m+1}{0}b^{m+1}v^{0}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\binom{m+1}{0}b^{m+1}u^{0}\,.$}
In the fifth line, we use the binomial theorem. $\Box$
We introduce improper Newton integrals of convergent formal power series.
@@
Let $u\in\mathbb{R}$ and $f(x)\in\mathbb{R}\{x\}$. If for every real sequence $(b_{n})$ going to $\infty$ the proper limit
@@
We call $L$ the improper Newton integral of $f(x)$ from $u$ to $\infty$.
@@
We prove the same three integral transformations as before.
@@
The third transformation is immediate as well.
@@
Let $f(x)\in\mathbb{R}\{x\}$ and $u,b\in\mathbb{R}$. The equality
@@
Proof. Let $f(x)$ and $u,b$ be as stated. Let $(c_{n})$ be any real sequence going to $\infty$. We assume that the integral $I$ on the left-hand side exists, and prove that the integral on the right-hand side exists too and equals $I$. Indeed,
@@
Similarly, if the integral on the right-hand side exists, then
[…]
$\Box$
We need to bound Newton integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)$ is in $\mathbb{Z}[x]$. We easily obtain the next bound based on two values of the primitive.
@@
$\text{$\sum_{n\geq 0}\frac{(-1)^{n}}{n!}x^{n+k}$ has the primitive $\sum_{n\geq 0}\frac{(-1)^{n}}{n!(n+k+1)}x^{n+k+1}$}\,.$}
Thus
$\big|\int_{0}^{i}x^{k}\cdot\mathrm{Exp}(-x)\big|\leq\sum_{n\geq 0}\frac{1}{n!}i^{n+k+1}+0=i^{k+1}\mathrm{e}^{i}\,.$}
$\Box$
It is crucial that the bound is exponential and not factorial in $k$.
The key step in Hilbert’s proof of the transcendence of $\mathrm{e}$ is Euler’s identity. We prove it in the language of improper Newton integrals of convergent formal power series.
Proposition 3.16
@@
$I_{k}=\int_{0}^{+\infty}x^{k}\cdot\mathrm{Exp}(-x)=\int_{0}^{+\infty}\sum_{n\geq k}\frac{(-1)^{n-k}}{(n-k)!}x^{n}=k!\,.$}
Proof. We proceed by induction on $k$. Let $(a_{n})$ be any real sequence going to $\infty$. Since
$\int\mathrm{Exp}(-x)=\sum_{n\geq 0}\frac{(-1)^{n}}{(n+1)!}x^{n+1}=1-\mathrm{Exp}(-x)\,$,}
we have
@@
Let $F_{k}=x^{k}\cdot\mathrm{Exp}(-x)$, $k\in\mathbb{N}_{0}$, and let $k\in\mathbb{N}$. By item 1 of Proposition 3.1 and the derivative of $\mathrm{Exp}(-x)$,
$F_{k}^{\prime}=\big(x^{k}\cdot\mathrm{Exp}(-x)\big)^{\prime}=kx^{k-1}\cdot\mathrm{Exp}(-x)-x^{k}\cdot\mathrm{Exp}(-x)=k\cdot F_{k-1}-F_{k}$
and $F_{k}=k\cdot F_{k-1}-F_{k}^{\prime}$. By item 4 of Proposition 3.1, $\int F_{k}^{\prime}=F_{k}-[x^{0}]F_{k}=F_{k}$. By Proposition 3.12 and induction,
@@
$=$ $k\cdot I_{k-1}-\lim_{n\to\infty}\big(F_{k}(a_{n})-F_{k}(0)\big)$
$=$ $k\cdot(k-1)!-0=k!\;.$
@@
where $a_{i}\in\mathbb{Z}$, $n\in\mathbb{N}$, and $a_{0}\neq 0$. We multiply this equality by the improper Newton integral
@@
where $p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}$ is as in Section 1. This integral exists by Propositions 3.12 and 3.16. Using Proposition 3.13 we get that
@@
for every $r\in\mathbb{N}$. The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as at the beginning, only now they are defined in an AB &\& SF way. We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded. The computation of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that now we calculate semiformally.
@@
Using Propositions 3.12 and 3.16, and the definition of the polynomial $p_{r}(x)$, we obtain the properties of $B_{r}$ as before.
It remains to bound, in an AB &\& SF way, the numbers
@@
has degree less than $(n+1)(r+1)$ and each coefficient in it is in absolute value at most $l^{r+1}$ where $l\in\mathbb{N}$ depends only on $n$. By Propositions 3.8 and 3.15 each of the above $n+1$ Newton integrals in $A_{r}$ is in absolute value at most
@@
Thus $|A_{r}|\leq c^{r}$ for every $r\in\mathbb{N}$ and some constant $c\geq 1$. We get a contradiction as before. $\Box$
The previous proof is semiformal and not formal due to the use of various semiformal operations and many values of formal power series. It is clearly an AB proof.